Fermi National Laboratory


Accelerator Update

Monday April 26
The day shift began with Operations monitoring store 3444 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter and MTest experiments taking beam.

At 10:15 AM, Operations halted stacking to investigate a problem with core emittances. Stacking resumed about twenty-five minutes later.

MCenter went into access at noon and came out at 1:30 PM.

MTest went into access at 3:16 PM and came out about four hours later.

A Main Injector (MI) expert began a study period at 3:21 PM.

Pbar stopped stacking and began a study period at 8:38 PM.

Tevatron (TeV) experts began a one-hour separator study at 9:36 PM.

Tuesday April 27
The midnight shift began with Operations monitoring store 3444 and the stash, with a Pbar expert conducting reverse proton studies, and with MiniBooNE, MCenter, and MTest taking beam.

At 1:41 AM, Operations reported a problem with Recycler emittances, which resulted in Recycler losing some of its stash.

Pbar completed their studies and resumed stacking at 4:41 AM.

At 6:43 AM, Operations reported more problems with Pbar emittances.

MCenter experimenters said at 8 AM that they'd be down and not taking beam for a few days.

Throughout the morning, MI experts began a 2.5MHz study at 9:16 AM, a slipstacking study at 10:13 AM, and then a 120 GeV orbit study at 11 AM.

TeV experts began an end-of-store study at noon and ended the study ten minutes later.

Operations terminated store 3444 at 12:50 PM.

A TeV expert began an abort study at 12:54 PM.

Operations made the Collider Detector at Fermilab (CDF) and D-zero (D0) ready to access their collision halls by 1:16 PM.

The Neutron Therapy Facility (NTF) began a study period at 2:30 PM.

All TeV accesses had ended by 2:32 PM.

At 3:11 PM, Booster/MI experts began cogging studies.

Pbar resumed stacking at 3:30 PM.

At 4:50 PM, Operations disabled MTest and put the TeV into shot setup.

At 6:35 PM, a TeV separator (B11H) sparked and kicked out the last bunch of protons. This caused emittances to blow up.

Operations established store 3446 with an initial luminosity of 13.44E30. But the beam problems were so bad neither colliding experiment could turn on their detectors. Pbar resumed stacking.

NTF finished their study at 7:09 PM.

Operators tried to smooth the beam with collimators at 7:38 PM. This allowed CDF to turn on, but D0 couldn't.

Operations terminated store 3446 at 8:16 PM.

TeV experts began a study period at 8:35 PM.

Operations reestablished beam to MTest at 10:16 PM.

Wednesday April 28
The midnight shift began with Operations monitoring the stash, with Pbar stacking, with TeV experts conducting studies, and with MiniBooNE and MTest taking beam.

At 1:16 AM, Operations put the TeV through a dry squeeze for the TeV studiers.

TeV studies ended around 3 AM.

A Linac RF station (LRF3) tripped off a number of times at 4:09 AM. It eventually reset itself. Then at 4:18 AM it did the same thing again.

Operations disabled the MTest beam and put the TeV into shot setup at 5 AM.

At 5:54 AM, the TeV sector A1 cryo system cold compressor tripped off. Operators aborted the protons in the TeV and informed the Cryo Coordinator.

Pbar resumed stacking at 6:25 AM.

Operators made D0 ready for an access at 6:53 AM.

Operations reestablished beam to MTest at 7:13 AM.

Plans for Today
The plans for today are to finish the A1 cold compressor work, to allow a TeV access for maintenance, and then establish a store later on today. Pbar will resume stacking and then shoot antiprotons to the Recycler.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/28/2004   email Fermilab